OpenStudy (anonymous):

Simplify the expression. Write the answer using scientific notation. 8(8.8 x 1012) A. 70.4 x 1012 B. 70.4 x 1024 C. 70.4 x 1013 D. 1.68 x 1013

OpenStudy (anonymous):

i know its not A is it C

OpenStudy (anonymous):

@hartnn

OpenStudy (raden):

is that 1012 or 10^12 ?

OpenStudy (raden):

i mean, for the original ur question

OpenStudy (unklerhaukus):

\[8\times (8.8 \times 10^{12})\\=(8\times 8.8 )\times 10^{12}\\=(64+6.4)\times10^{12}\\=70.4\times10^{12}\\=7.04\times10^{12+1}\]

OpenStudy (unklerhaukus):

or you sure those options are right? because the mathematically equivalent option is Not in scientific notation

OpenStudy (raden):

it is E : there is no options be right :)

OpenStudy (anonymous):

its A